OpenStudy (anonymous):

Can you do the integral test on an increasing function when analysing an increasing series?

OpenStudy (anonymous):

Or, more generally, what restrictions are there as to *when* you can use the integral \[\int\limits_?^{\infty} f(x)\] to see whether \[ \sum_{k=0}^{\infty}f(k)\] converges?